### v1

# A Comparative Analysis of Large Language Models for Code Documentation Generation

Abstract.
This paper presents a comprehensive comparative analysis of Large Language Models (LLMs) for generation of code documentation. Code documentation is an essential part of the software writing process. The paper evaluates models such as GPT-3.5, GPT-4, Bard, Llama2, and Starchat on various parameters like Accuracy, Completeness, Relevance, Understandability, Readability and Time Taken for different levels of code documentation. Our evaluation employs a checklist-based system to minimize subjectivity, providing a more objective assessment. We find that, barring Starchat, all LLMs consistently outperform the original documentation. Notably, closed-source models GPT-3.5, GPT-4, and Bard exhibit superior performance across various parameters compared to open-source/source-available LLMs, namely LLama 2 and StarChat. Considering the time taken for generation, GPT-4 demonstrated the longest duration, followed by Llama2, Bard, with ChatGPT and Starchat having comparable generation times. Additionally, file level documentation had a considerably worse performance across all parameters (except for time taken) as compared to inline and function level documentation.
Code documentation, Large Language Models
conference: 46th International Conference on Software Engineering; April 2024; Lisbon, Portugalccs: General and reference Evaluationccs: Computing methodologies Natural language generationccs: Software and its engineering Documentation
1. Introduction
In the field of artificial intelligence, Large Language Models (LLMs) are revolutionizing how we handle information and develop software (Ross et al., 2023). These models, built on massive data sets and sophisticated algorithms, have moved beyond mere convenience to become vital in the software creation process, offering unprecedented assistance in code generation, debugging, and indeed, code documentation. As software grows more complex and development speeds up, our dependence on these digital assistants is increasing. This makes it crucial to regularly assess their effectiveness and accuracy.
Code documentation generation using LLMs (Rai et al., 2022), (Ahmed and Devanbu, 2023) is essentially derived from the root concept of code interpretability by an LLM and then adds on layers to it. To generate documentation, the LLM must be able to understand the code, explain the code, and then finally format the code into an easily understandable, legible and usable form.
High-quality code documentation is extremely important (Lethbridge et al., 2003). It is the foundation of good maintenance, growth, and teamwork in software development. Good documentation makes code easy to use and understand. This helps teams work together smoothly and keeps software projects healthy for a long time. Therefore, we need to pay close attention to the quality of documentation made by Large Language Models (LLMs) as it has a direct impact on how well and successfully human developers can use it.
LLMs have been studied in terms of code explanation and code summarizing, but this has mostly been done from an educative perspective (MacNeil et al., 2022), (Sarsa et al., 2022), (Leinonen et al., 2023). Moreover, not a lot of research has been done comparing the performance of closed source and open source LLMs for code documentation tasks (Alizadeh et al., 2023), (Gudibande et al., 2023). This study aims to analyze the effectiveness of using both open source and closed source LLMs to generate industrial level code documentation.
With this paper, we aim to explore the following objectives:
  • Compare leading LLMs on their ability to generate various levels of code documentations.
  • Compare the performances of LLMs with the human generated documentation for the same code base.
  • Compare the performances of closed/private sourced LLMs with LLMs whose sources are publicly available.
In doing so, we aim to pave the path for future research and development, pinpointing areas where LLMs excel and identifying gaps that require attention in the realm of code explanation and documentation.
2. Related Work
Code documentation generation has been a subject of considerable research, with previous studies falling into two broad categories.
  • The first category encompasses efforts that share a similar conceptual foundation but were hampered by technological limitations of their time. Early efforts in this space were limited by the capabilities of the available tools and frameworks, hindering the full exploration of the ideas (Che, [n. d.]), (Meneses, [n. d.]), (Khan and Uddin, 2022).
  • The second category of research has focused on code explanation, summarization, and interpretation, primarily driven by educational objectives. While these studies have made significant strides in enhancing code understanding using LLMs, they may not be directly translatable to the specific needs of industrial-level documentation generation. Educative perspectives often prioritize clarity and teaching over formatting and detailed context necessary for comprehensive industry-level documentation (MacNeil et al., 2022), (Leinonen et al., 2023), (Sarsa et al., 2022), (Su et al., 2023), (Naveed et al., 2023).
Additionally, the studies comparing source-available LLMs with close-source proprietary LLMs are limited to general natural language tasks and overall performances and very little work has been done to compare their performances on code specific tasks and specifically code documentation generation (Alizadeh et al., 2023), (Gudibande et al., 2023).
Distinct from the above mentioned categories, the present study addresses the challenges associated with the generation of industry-level code documentation. Unlike past efforts constrained by technological limitations, the current investigation utilizes the advancements in LLMs to explore the full potential of code documentation generation. Our work differentiates itself from research centered on code summarization for educational purposes by explicitly targeting the additional layer of formatting required in industry-level documentation, acknowledging the need for precise and well-structured information in real-world software development scenarios. Moreover, we also compare the performance between source-available LLMs and closed source proprietary LLMs from a documentation generation and general code interpretation perspective.
3. Methodology
3.1. Levels of documentation
To conduct a thorough analysis of the LLMs’ capabilities in generating code documentation, we decided to categorise the documentations into four distinct categories/levels (Rai et al., 2022):
  • Inline level documentation
  • Function level documentation
  • Class level documentation
  • Folder level documentation
Each level represents a different scope and complexity of documentation, providing insights into how well each LLM handles varying degrees of detail and abstraction in technical writing.
3.1.1. Inline Level:
It involves the generation of comments within the code, explaining specific lines or blocks of code. Inline documentation is crucial for clarifying complex or non-obvious parts of the code, enhancing readability and maintainability. It can also be very helpful in explaining a piece of code line by line to new coders. This will help evaluate the LLMs’ ability to produce concise, relevant, and helpful inline comments that enhance understanding without cluttering the code.
3.1.2. Function level:
At the function level, the focus shifts to documenting individual functions or methods. Good function-level documentation explains the purpose, parameters, return values, exceptions, and usage examples of functions. This assessment will measure how well each LLM can summarise the functionality and provide clear, accurate descriptions for function usage.
3.1.3. Class level:
Class level documentation provides an overview of a class, its purpose, and its interactions with other classes. It often includes descriptions of key attributes and methods within the class. This level of documentation is critical for understanding the design and architecture of the software.
3.1.4. Folder Level:
Folder Level Documentation is the highest level of documentation involving creating overviews for entire folders or modules, summarising the functionalities and interactions of multiple classes or files within them. We did not perform the analysis on folder level documentation because of technological limitations in most of the LLMs (except for GPT-4) which could not handle multiple file inputs at the same time.
3.2. Evaluation parameters and metrics
We carefully picked evaluation criteria and their measures to thoroughly analyse the database. Our goal was to cover all aspects of documentation, making sure we consider everything. Importantly, we wanted our assessment to be as objective as possible. So, the metrics we chose are designed to give a fair and straightforward evaluation of the documentation. This way, we aim to capture the full picture of documentation quality without relying too much on personal opinions. The evaluation parameters and their corresponding metrics are listed down below:
  1. (1)
    Accuracy:
    Accuracy is a check of correctness of the documentation generated. It checks if the documentation generated correctly describes the code or not. The accuracy is rated on a scale of 1-3, with 1 being the least accurate and 3 being perfectly accurate.
  2. (2)
    Completeness:
    Completeness of the documentation indicated the extent to which all the (important) parts of the code have been covered or not. Completeness is rated using a 5 point checklist on a scale of 0-5, on the basis of the number of checkpoints the documentation satisfied. Following are the checkpoints for the various documentation levels:
    • For file level: Brief description of the file/script’s purpose is included, List of dependencies is provided, Instructions or notes on how to use the script are included, List of main functions or classes defined in the file (if applicable), Any special notes or considerations relevant to the script (e.g., context of use, limitations).
    • For function level: The description of the function’s purpose is clear, All parameters are described with type and purpose, Return values are explained(if any), Exceptions or errors the function may raise are documented.(if any), Any additional notes or warnings are included.
    • For inline level: Complex code blocks have accompanying comments, Non-obvious algorithmic decisions are explained (if any), Workarounds or technical debt is noted (if any), Any assumptions in the code are stated, Deprecated methods or upcoming changes are flagged. (depends on LLM)
  3. (3)
    Relevance: Relevance shows how relevant the generated documentation is to the actual code, essentially it’s ability to stick to the subject matter of the code and not go off topic. Relevance is rated on a scale of 1 to 4 as explained below:
    1 - No relevant information; completely off-topic or non-existent.
    2 - Somewhat relevant; mixes essential and non-essential information.
    3 - Mostly relevant; includes most of the key details necessary for understanding the function.
    4 - Fully relevant; every piece of information helps understanding the function.
  4. (4)
    Understandability: Understandability indicates how well a person reading the documentation is able to understand what the piece of code means/ how to use the code. Rating understandability is very much dependent on the level of experience and skill a user has in software development. Hence, there is a scope of understandability being a little subjective as an evaluation parameter, and is evaluated on the scale of 1 to 4, where each value means the following:
    1 - Completely unintelligible; impossible to understand.
    2 - Somewhat understandable; could benefit from clearer language or examples.
    3 - Mostly understandable; fairly easy to grasp but could be improved.
    4 - Fully understandable; extremely clear and easy to understand, even for someone unfamiliar with the topic.
  5. (5)
    Readability: Readability rates the formatting of the documentation. A well formatted documentation makes it convenient for the reader to go through the documentation. Readability is rated using a 5-point checklist on a scale of 0-5, on the basis of the number of checkpoints the documentation satisfies. The checkpoints are given below:
    • File level: Header comment has clear delineation (e.g., lines or asterisks). (Documentation is clearly separated from the code), Consistent indentation is used for nested information, Lines are properly aligned(horizontal scrolling is absent), Is metadata like author, date, etc. present, Presence of spacing between Sections.
    • Function level: Is the docstring clearly distinguishable from the code,, Is there consistent formatting in the docstring (e.g., for parameters, returns), Any examples within the docstring are properly indented, Line breaks are used to separate different sections within the docstring, Is the docstring structured logically (e.g., description, parameters, returns, examples)?
    • Inline level: Are comments placed close to the code they describe or do comments follow the same indentation as the code they are describing, Comments are brief and do not exceed the length of the code line, Consistent style is used for single-line and multi-line comments, Is there adequate spacing around comments for clarity, Are comments placed logically in the code for easy association with the relevant code?
  6. (6)
    Time: This is a simple measure of the time taken by the model to fully generate the response after the prompt has been entered.
3.3. Prompts
The prompts used as inputs to the LLMs were carefully constructed using the concepts of prompt engineering to generate the best possible responses from the large language models. The same prompt was used across models for the same level of documentation to maintain homogeneity and fairness while comparing the performances of models. The prompts were generally structured in a way such that they included the following:
  • Persona: The LLM was assigned a persona
  • Context: A context was given to the LLM under which the documentation is to be generated
  • Task: The LLM was given the task using clear action words.
  • Format: The LLM was specified the format in which the documentation was to be generated.
The prompts used for each level of documentation were slightly different, and are given below:
File level: As a code documentation assistant, you are responsible for documenting at the file/script level of the given code snippet. When provided a file level code, your approach involves adding a header comment at the top of the file. This comment should be the documentation for the code and include all relevant information needed to understand or use the script. Code is as follows: ¡insert code¿
Function level: As a code documentation assistant, you are programmed to document at the function level of the given code snippet. Your approach involves placing comments directly under the def statement of the function. The output should be the entire code along with the documentation of the function written. Code is as follows: ¡insert code¿
Inline level: As a code documentation assistant, you are assigned to document at the in-line level of the given code snippet. When in-line comments are needed, you insert comments within the code itself. The output should be the entire code, along with the documentation you’ve added. Code is as follows: ¡insert code¿
3.4. LLMs Used
In our comparative analysis of code documentation generation capabilities among various large language models (LLMs), both open-source/source-available and closed-source models were utilized. The LLMs included in our study are GPT-3.5, GPT-4, Bard, LLama2, and Starchat.
The closed source proprietary LLMs used are the following:
  • GPT-3.5:(Brown et al., 2020) Developed by OpenAI, GPT-3.5 is an advanced iteration of the Generative Pre-trained Transformer models, boasting around 175 billion parameters. It is one of the most popular LLMs used as of right now.
  • GPT-4:(OpenAI, 2023) Also from OpenAI, GPT-4 is a more sophisticated and larger model than its predecessor, with improved performance in nuanced language understanding and generation. It is built on 1.76 trillion parameters and is also able to handle visual inputs along with textual ones.
  • Bard:(Manyika and Hsiao, [n. d.]) Bard is a closed-source LLM developed by Google and is built on 137 billion parameters. It is one of the major players in the LLM space after OpenAIs chatGPT.
The open source/source available LLMs used are the following:
  • LLama2:(Touvron et al., 2023) LLama2 is a source-available LLM developed by Meta. It has different model sizes with 7, 13 and 70 billion parameters. The one used for this study has 70 billion parameters.
  • Starchat:(Li et al., 2023) Starchat 2 is a fine tuned version of the StarCoder code completion focused model and is one of the more commonly used code focused open source LLM. It is built on around 15 billion parameters.
3.5. Data Analysis
A set of 14 python code datasets were collected from various sources (like github and other open source repositories and websites). The codebases from which these code snippets were collected are all well reputed and very popular and have a very high level and well curated documentation. These sources included scikit-learn, bitcoin codebase etc.
These code snippets are were then individually run through all the LLMs that we had collected and were evaluated on each of the parameter that was mentioned above and the assessed data was then collected for evaluation.
4. Evaluation
We conducted a thorough analysis of 14 Python code datasets, each accompanied by its respective industrial standards documentation for reference and comparison. We then ran various forms of analyses over the dataset, following are the results and observations:
4.1. Overall Analysis
Table 1 and 2 present a summary of the descriptive statistics like mean, standard deviation, range and distribution of values etc. for various parameters, including Accuracy, Completeness, Relevance, Understandability, Readability, and Time Taken.
Metric Accuracy Completeness Relevance
count 84.000 84.000 84.000
mean 2.860 4.192 3.833
std 0.445 1.206 0.567
min 0.000 0.000 0.000
25% 3.000 4.000 4.000
50% 3.000 5.000 4.000
75% 3.000 5.000 4.000
max 3.000 5.000 4.000
Table 1. Accuracy, Completeness and Relevance averaged across all LLMs
Metric Understandability Readability Time Taken
count 84.000 84.000 84.000
mean 3.551 4.525 16.977
std 0.766 0.935 17.994
min 0.000 0.000 0.000
25% 3.000 4.000 0.000
50% 4.000 5.000 11.000
75% 4.000 5.000 30.250
max 4.000 5.000 75.000
Table 2. Understandability, Readability and Time Taken averaged across all LLMs
4.2. Comparison across different LLMs
To enhance the clarity of our findings, we utilized bar graph visualizations to compare mean ratings across different language models for each parameter.
Figure 1. Model wise accuracy comparison
In figure 1, we can see that there is a significant drop in average accuracy in StarChat, meanwhile the rest of the LLMs have almost on par accuracy as compared to the original documentation.
Figure 2. Model wise completeness comparison
In figure 2, we can see that GPT-3.5 and GPT significantly outperform even the original documentation in terms of completeness and thoroughness of the code documentation, meanwhile Bard and LLama 2 have somewhat the same, and slightly better and StarChar has slightly worse performance in terms of completeness than the original documentation.
Figure 3. Model wise relevance comparison
In figure 3, LLama 2 and StarChat have a slightly worse performance as compared to GPT-3.5 and 4 and Bard, which are on par with the original documentation.
Figure 4. Model wise understandability comparison
From figure 4, we can see that the documentation generated by GPT-3, GPT-4 and Bard have better understandability as compared to the original documentation. LLama 2 and StarChat have a somewhat similar performance as the original documentation.
Figure 5. Model wise readability comparison
From figure 5, we can say that all the models have high Readability, and most of them outperform the original documentation by small margins.
Figure 6. Model wise Time taken comparison
From figure 6, we can say that the time taken by GPT-4 to generate the documentation is very high as compared to the other models. GPT-3.5 and StarChat generate the documentation in the shortest amount of time. Bard and LLama 2 lie somewhere in between. This is the only criteria observed where GPT-4 has performed the worst by a large margin.
4.3. Correlation between different metrics
Figure 7. Metrics correlation heatmap
As we can see from the correlation heatmap in figure 7, there is a somewhat strong correlation between Accuracy, Understandability and Relevance (each about 0.7-0.75). Other than this, there is not strong correlation between the other parameters. It is also interesting to note that the time taken by each LLM is completely independent of all the other parameters.
Figure 8. Documentation level wise metrics comparison
Figure 9. Documentation level wise time taken comparison
From figure 8, we can see that performance of the LLMs while generating file level documentation has a consistently worse performance as compared to the performance while generating inline and function level documentation in all parameters.
From figure 9, it can be seen that even though file level documentation generated is rated less, it still takes the least amount of time and documentation is generated faster than the other 2 levels of documentation.
4.4. Statistical Analysis
For statistical analysis, ANOVA (Analysis of Variance) (St»hle and Wold, 1989) is employed to assess the significance of differences among group means in multi-group comparisons. This method is particularly essential when evaluating multiple models across various parameters, such as performance metrics. ANOVA utilizes F-statistics and p-values to determine whether observed variances across models are statistically significant or occur by chance, thereby providing a rigorous and technical approach to model evaluation and comparison.
Criterion F-statistic P-value
Accuracy 1.1803 0.3271
Completeness 3.9874 0.0030
Relevance 3.1599 0.0123
Understandability 2.2788 0.05562
Readability 1.0665 0.3861
Time Taken 29.5365 $1.0497\times 10^{-16}$
Table 3. Statistical Analysis Results
  • Accuracy:
    The F-statistic is relatively low, and the p-value is higher than the common alpha level of 0.05. This suggests that there is not enough statistical evidence to conclude that the model significantly affects accuracy.
  • Completeness:
    The F-statistic is higher, and the p-value is below 0.05. This indicates a statistically significant effect of the model on completeness.
  • Relevance:
    A higher F-statistic and a p-value below 0.05 suggest a significant effect of the model on relevance.
  • Understandability:
    The F-statistic indicates some effect, but the p-value is slightly above 0.05. This suggests that the model’s impact on understandability is borderline significant; it’s close to being significant but doesn’t quite reach the conventional threshold.
  • Readability:
    Both the F-statistic and the p-value indicate that the model does not have a significant effect on readability.
  • Time Taken:
    The very high F-statistic combined with an extremely low p-value (practically zero) strongly suggests that the model has a significant effect on the time taken.
5. Conclusion
Following are the findings of our comparative analysis:
  • With the exception of Starchat, all Large Language Models (LLMs) exhibit either equal or superior performance when compared to the original documentation. Starchat consistently yielded suboptimal results when compared to the original documentation.
  • Closed-source models, including GPT-3.5, GPT-4, and Bard, consistently outperform their open-source counterparts, Llama2 and Starchat, across a majority of parameters in our evaluation rubric. Notably, when considering the time taken for generation, GPT-4 demonstrated the longest duration, followed by Llama2, Bard, with ChatGPT and Starchat having comparable generation times.
  • Additionally, file level documentation had a considerably worse performance across all parameters (except for time taken) as compared to inline and function level documentation.